OpenStudy (anonymous):

$970 is invested for one year at 14% simple interest is earned.....the simple interest is ?

OpenStudy (anonymous):

Please use the formula I suggested here: http://openstudy.com/study#/updates/4f4b35dee4b0acf2d9fd3f59

OpenStudy (anonymous):

SI= PRT/100 = 970 * 14 /100 = 9.7 * 14 = $ 135.8